OpenStudy (anonymous):

For questions 3-4 solve the equation using square roots. A. x^2-81=0 B. 3x^2-147=0

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

is A -9,9 and B no real number solutions

OpenStudy (anonymous):

@phi @thomaster

OpenStudy (phi):

for A, add 81 to both sides. you get x^2 = 81 take the sqr of both sides: x= ±sqr(81)= ±9

OpenStudy (phi):

for B, first divide both sides (and all terms) by 3 to get x^2- 49=0 now solve the same as in A

OpenStudy (anonymous):

\[-\sqrt{9} \sqrt{9}\] is that right for A

OpenStudy (anonymous):

and b -49 49

OpenStudy (phi):

for A, it is \(\sqrt{81} \) which simplifies to +9 or -9

OpenStudy (anonymous):

oh so i switched the numbers ok

OpenStudy (phi):

for B x^2- 49=0 add 49 to both sides x^2 -49 + 49= 0+49 simplify (-49+49 on the left is 0) x^2 = 49 now take the square root of both sides

OpenStudy (anonymous):

7, -7 is the square root of 49, -49

OpenStudy (phi):

yes

OpenStudy (phi):

notice you run into trouble if the problem was x^2 + 49 =0 we would add -49 to both sides: x^2 + 49 - 49 = 0 - 49 simplify to x^2 = -49 <-- TROUBLE sqr of a negative number